# On the Capacity of Self-Attention

### v1

@@ 3 Modeling the Self-Attention Mechanism
We model the key–query (QK) channel of a single self‑attention layer for RGR, retaining permutation symmetry and parameter sharing while omitting components that do not affect the binary edge decision (softmax, $1/\sqrt{d_{k}}$ scaling, and values; see Appendix. C). The input is an ordered context of distinct vertices of length $\ell\!\leq\!m$, given as $\mathcal{C}=(v_{i_{1}},\dots,v_{i_{\ell}})$, one each to $\ell$ attention units. Each $v\in V$ is described using a unique embedding $\mathbf{x}_{v}\in\mathbb{R}^{d_{\text{model}}}$. Positional information is not explicitly modeled; if needed, positions can be incorporated by treating $(\text{token},\text{position})$ as distinct vertices.
@@
Each attention unit with a single head uses the same shared projection matrices $W_{Q},W_{K}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$. For each $v_{i_{p}}\!\in\!\mathcal{C}$,
$\mathbf{q}_{i_{p}}=\mathbf{x}_{i_{p}}W_{Q},\qquad\mathbf{k}_{i_{p}}=\mathbf{x}_{i_{p}}W_{K}.$
The unnormalized score from source $v_{i_{p}}$ to target $v_{i_{q}}$ is
$S_{pq}=\mathbf{q}_{i_{p}}\cdot\mathbf{k}_{i_{q}}^{\top}.$
We declare an edge $(v_{i_{p}},v_{i_{q}})$ present iff $S_{pq}>\tau$ for a global threshold $\tau$. Only pairs inside $\mathcal{C}$ are tested.
Multi‑head extension.
With $h$ heads, each head $k$ has $(W_{Q}^{(k)},W_{K}^{(k)})\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ and produces $S_{pq}^{(k)}$. We aggregate per pair by
$S_{pq}^{\max}=\max_{k\in\{1,\dots,h\}}S_{pq}^{(k)},\quad\text{and decide}\quad(v_{i_{p}},v_{i_{q}})\in E\ \Leftrightarrow\ S_{pq}^{\max}>\tau.$
This “OR‑of‑heads” view matches the common specialization picture. Also, replacing $\max$ by log‑sum‑exp yields an equivalent classifier after a global threshold shift; see App. C.
@@
A construction for RGR maps a graph $G=(V,E)$ to weights $\{(W_{Q}^{(k)},W_{K}^{(k)})\}_{k=1}^{h}$ and a threshold $\tau$ that realize the correct edge decisions for all contexts $\mathcal{C}$, regardless of length $\ell$. We measure complexity by the total key dimension
$D_{K}\;=\;h\,d_{k},$
since QK is implemented as two batched multiplications with concatenated weights
$W_{Q}^{\text{cat}}=[W_{Q}^{(1)}|\cdots|W_{Q}^{(h)}]\in\mathbb{R}^{d_{\text{model}}\times(hd_{k})}\text{ and }W_{K}^{\text{cat}}=[W_{K}^{(1)}|\cdots|W_{K}^{(h)}]\in\mathbb{R}^{d_{\text{model}}\times(hd_{k})};$
the dominant cost and parameter footprint scale with $h\,d_{k}$ rather than $h$ or $d_{k}$ alone (see Appendix C). [^2] All statements remain (up to a factor of two) if one counts $D_{Q}+D_{K}$. Our goal is to minimize $D_{K}$ over a graph family $\mathcal{G}_{m,m^{\prime}}$ for a given $d_{\text{model}}$.
In Appendix C, we provide further justification for many of our model choices. Specifically, we address the use of thresholding instead of scaling/softmax, our head aggregation rule, as well as the lack of a value pathway.
4 Explicit Constructions for Relational Graph Recognition
In this section, we provide explicit constructions for the attention weights that solve RGR within the idealized model of Section 3. These constructions establish an upper bound on the $D_{K}$ required to solve RGR, providing a concrete measure of the self-attention mechanism’s capacity for this task. Each design is proven to perform with high probability (w.h.p.), meaning at least $1-m^{-\gamma}$ for some constant $\gamma>2$. We begin with a brief warm‑up sketch—recognizing permutation graphs with one‑hot embeddings—then move to compressive embeddings. We then generalize to arbitrary graphs. Throughout, $i$ indexes the source and $j$ a candidate target in $E$; keys are tied to targets and queries are tied to sources.
Construction I: Permutation Graphs with One‑Hot Embeddings (Warm-up/sketch).
We consider a permutation graph $G$ on $m$ items with edges $E=\{(v_{i},v_{\pi(i)})\}$ and one‑hot node encodings $\mathbf{x}_{i}=\mathbf{e}_{i}\in\mathbb{R}^{m}$ (so $d_{\text{model}}=m$). A single head ($h=1$) suffices. Assign each target $v_{j}$ a random binary signature as its key $\mathbf{k}_{j}$ by drawing $W_{K}\in\{0,1\}^{m\times d_{k}}$ with i.i.d. $\mathrm{Bernoulli}(p)$ entries (e.g., $p=1/4$) and setting $\mathbf{k}_{j}=\mathbf{e}_{j}W_{K}$. For each source $v_{i}$, set the query to the signature of its target: $\mathbf{q}_{i}=\mathbf{k}_{\pi(i)}$, i.e., the ii‑th row of $W_{Q}$ equals the $\pi(i)$‑th row of $W_{K}$. With $d_{k}=C\log m$ and threshold $\tau=\tfrac{p+p^{2}}{2}\,d_{k}$, one has
$S_{i,\pi(i)}=\mathbf{k}_{\pi(i)}\!\cdot\!\mathbf{k}_{\pi(i)}\sim\mathrm{Binomial}(d_{k},p),\qquad S_{ij}=\mathbf{k}_{\pi(i)}\!\cdot\!\mathbf{k}_{j}\sim\mathrm{Binomial}(d_{k},p^{2})\;\ (j\!\neq\!\pi(i)).$
Chernoff bounds and a union bound over all $(i,j)$ yield simultaneous separation $S_{i,\pi(i)}>\tau>S_{ij}$ w.h.p. when $d_{k}=\Theta(\log m)$, so a single head recovers all edges. This matches the $\Omega(\log m)$ lower bound (Section 7) and the same argument extends to softmax scoring (Appendix C). Full details—algorithm and proof—are deferred to Appendix B as a “warm‑up” proof; Construction II below generalizes this scheme to compressive embeddings, but first we point out that a separation of the the type described is sufficient to handle contexts of any length.
Lemma 4.1 (Context‑robustness).
Fix parameters $\{(W_{Q}^{(k)},W_{K}^{(k)})\}_{k=1}^{h}$ and a threshold $\tau$. Let the aggregated score be $S_{ij}^{\max}:=\max_{k}S_{ij}^{(k)}$. If, simultaneously for all $i\in V$,
$S_{i,\pi(i)}^{\max}>\tau\quad\text{and}\quad S_{ij}^{\max}<\tau\ \ \text{for all }j\neq\pi(i),$
then for every context $\mathcal{C}\subseteq V$ and every source $i\in\mathcal{C}$: (i) if $\pi(i)\in\mathcal{C}$ then $S_{i,\pi(i)}^{\max}>\tau$ and $S_{ij}^{\max}<\tau$ for all $j\in\mathcal{C}\!\setminus\!\{\pi(i)\}$; (ii) if $\pi(i)\notin\mathcal{C}$ then $S_{ij}^{\max}<\tau$ for all $j\in\mathcal{C}$. Hence the same parameters recognize $E|_{\mathcal{C}}$ for every $\mathcal{C}$ and every length $\ell$.
Proof.
Restricting from $V$ to $\mathcal{C}$ only removes candidate targets. If $\pi(i)\in\mathcal{C}$, the inequalities in $(\star)$ remain true after removing all $j\notin\mathcal{C}$. If $\pi(i)\notin\mathcal{C}$, every remaining $j\in\mathcal{C}$ satisfies $j\neq\pi(i)$, so $S_{ij}^{\max}<\tau$ by $(\star)$. ∎
@@ 4 Explicit Constructions for Relational Graph Recognition
We now extend the permutation case to the compressive regime $d_{\text{model}}\ll m$ under a Gaussian unit‑norm embedding. [^3] Random spherical codes are nearly orthogonal—inner products concentrate tightly around 0—which lets us obtain clean dot‑product thresholds and $O(\log m)$ scaling in our separation arguments. This cosine‑geometry is also standard and effective in practice: many systems explicitly constrain features to a hypersphere (e.g., NormFace and ArcFace in face recognition; Spherical Text Embedding in NLP; spherical objectives in metric learning). See [62, 13] for concentration/JL background and [66, 14, 44, 73] for representative uses of unit‑sphere embeddings. Each item $v_{i}$ is embedded as a fixed vector $\mathbf{x}_{i}\in\mathbb{R}^{d_{\text{model}}}$ drawn i.i.d. as $\tilde{\mathbf{x}}_{i}\sim\mathcal{N}(0,I/d_{\text{model}})$ and then $L_{2}$‑normalized, i.e., $\mathbf{x}_{i}=\tilde{\mathbf{x}}_{i}/\|\tilde{\mathbf{x}}_{i}\|_{2}$. Write $X\in\mathbb{R}^{m\times d_{\text{model}}}$ for the matrix with ii‑th row $\mathbf{x}_{i}^{\top}$. Given such an embedding and permutation $\pi$, our goal is to construct attention parameters that recognize $G$.
Multi‑Head Algorithmic Construction
The fundamental challenge with embeddings is that the input $\mathbf{x}_{i}$ is a superposed representation of the node’s identity. Our construction first approximately inverts the embedding process, projecting the $d_{\text{model}}$-dimensional vector $\mathbf{x}_{i}$ back into the $m$-dimensional one-hot space using the transpose of the embedding matrix. We then apply the logic from the one-hot case. However, doing this with a single head yields too much noise due to the inversion being only approximate. We mitigate this noise by using multiple attention heads, where each is responsible for recognizing the outgoing edges from a disjoint subset of sources. This results in smaller individual heads, and thus less noise.
Algorithm 1 Construction for Permutation Graphs with Compressive Embeddings
@@
2:Parameters: Number of heads $h=\frac{m}{d_{\text{model}}}$; per‑head key/query dimension $d_{k}=C\log m$ for a sufficiently large absolute constant $C$. Set Threshold: $\tau=\frac{1}{2}d_{k}$.
3:Partition sources and targets: Split $V$ into $h$ disjoint blocks $V_{1},\dots,V_{h}$ of size $|V_{k}|=d_{\text{model}}$. For each head $k$, define its target set $T_{k}\;:=\;\{\pi(s):s\in V_{k}\},$ which is a permutation of $V_{k}$. Head $k$ is responsible for sources in $V_{k}$ and targets in $T_{k}$ (a permutation of $V_{k}$).
4:Random signatures: Draw $W_{\mathrm{sig}}\in\{\pm 1\}^{m\times d_{k}}$ with i.i.d. Rademacher entries; let $\mathbf{w}_{j}$ be its jj‑th row.
5:Ideal one‑hot‑space templates (for each head $k$):
6:   $\textbf{Query Matrix:}\;W^{\prime}_{Q,(k)}\in\mathbb{R}^{m\times d_{k}}$ with row $i$ equal to $\mathbf{w}_{\pi(i)}$ if $i\in V_{k}$, and $\mathbf{0}$ otherwise.
7:   $\textbf{Key Matrix:}\;W^{\prime}_{K,(k)}\in\mathbb{R}^{m\times d_{k}}$ with row $j$ equal to $\mathbf{w}_{j}$ if $j\in T_{k}$, and $\mathbf{0}$ otherwise.
8:Project back to model space (approximate de‑embedding):
@@
Theorem 4.2 (Multi‑head recognition under Gaussian unit‑norm embeddings).
@@
Assume the setup above with $h=\frac{m}{d_{\mathrm{model}}}$ heads, per‑head dimension $d_{k}=C\log m$ for a sufficiently large absolute constant $C$, and threshold $\tau=\tfrac{1}{2}d_{k}$. If $d_{\mathrm{model}}\geq c_{0}\log m$ for a sufficiently large absolute constant $c_{0}$, then with probability at least $1-m^{-3}$ over the draw of $(X,\text{signatures})$,
$\forall i\in V\;\exists\,k\in[h]\text{ with }i\in V_{k}:\quad S^{(k)}_{i,\pi(i)}>\tau\quad\text{and}\quad S^{(k)}_{ij}<\tau\ \ \forall j\neq\pi(i).$
@@
Consequently, max‑pooling over heads correctly recognizes all edges and $D_{K}=h\,d_{k}=\Theta\!\Big(\frac{m\log m}{d_{\mathrm{model}}}\Big).$
Proof sketch (intuition).
Write $G=XX^{\top}$ and note $G\approx I$ w.h.p. when $d_{\text{model}}\gtrsim\log m$. For a source $i\in V_{k}$, the de‑embedded vector $\mathbf{u}_{i}:=\mathbf{x}_{i}X^{\top}$ equals $\mathbf{e}_{i}+\boldsymbol{\delta}_{i}$ where the leakage $\boldsymbol{\delta}_{i}$ has small $\ell_{2}$ mass on the head‑local blocks $V_{k}$ and $T_{k}$. Using Rademacher signatures, the score decomposes as
$S^{(k)}_{ij}=\underbrace{\mathbf{w}_{\pi(i)}\!\cdot\!\mathbf{w}_{j}\cdot\mathbb{I}(j\in T_{k})}_{\text{Signal}}\ +\ \underbrace{\text{terms linear/quadratic in }\boldsymbol{\delta}_{i},\boldsymbol{\delta}_{j}}_{\text{Noise}}.$
At the true edge $j=\pi(i)$, Signal equals $d_{k}$ exactly; at non‑edges it is 0 (if $j\notin T_{k}$) or a sub‑Gaussian fluctuation $O(\sqrt{d_{k}\log m})$ (if $j\in T_{k}\setminus\{\pi(i)\}$). Concentration results for random spherical codes implies (i) per‑head leakage energy $\|\boldsymbol{\delta}_{i,S}\|_{2}^{2}\lesssim 1$ for $S\in\{V_{k}\!\setminus\!\{i\},T_{k}\}$ and (ii) small cross‑terms $\langle\boldsymbol{\delta}_{i,\pi^{-1}(T_{k})},\boldsymbol{\delta}_{j,T_{k}}\rangle\lesssim\sqrt{\frac{\log m}{d_{\text{model}}}}$. Thus each Noise component concentrates to $O(\sqrt{d_{k}\log m})+o(d_{k})$; choosing $d_{k}=C\log m$ and $d_{\text{model}}\geq c_{0}\log m$ makes Noise $\tfrac{1}{4}d_{k}$, while Signal at the true edge is $d_{k}$. Setting $\tau=\tfrac{1}{2}d_{k}$ yields $S^{(k)}_{i,\pi(i)}>\tau>S^{(k)}_{ij}$ uniformly, and a union bound over $(i,j,k)$ gives the stated probability. A full proof with explicit constants and concentration lemmas appears in Appendix B.
Consequences.
The total key budget satisfies $D_{K}=O\!\big(\frac{m\log m}{d_{\text{model}}}\big)$, which matches our lower bound up to constants. Moreover, the non‑edge bounds hold head‑wise, so for any $j\neq\pi(i)$ we have $S_{ij}^{(k)}<\tau$ for all $k$ and hence $S_{ij}^{\max}<\tau$, while $S_{i,\pi(i)}^{\max}>\tau$ at the true target. By Lemma 4.1, the same parameters correctly recognize $E|_{\mathcal{C}}$ for every context $\mathcal{C}$ and for every context length. (As in Construction I, the thresholding analysis translates to softmax; see Section C.)
Construction III: More General Embeddings (Summary).
In Appendix B, we abstract the geometric assumptions used in Construction II into a reusable, block‑level condition on the embedding $X$: restricted self‑incoherence (Def. B.4), which requires the existance of a block size parameter $B$ such that (a) after rescaling by a global factor $\mu$, the Gram matrix $XX^{\top}$ has diagonals close to $1$; (b) when a row of $XX^{\top}$ is restricted to any subset of at most $B$ indices, the off‑diagonal “leakage” energy is small (so projecting $\mathbf{x}_{i}$ back with $X_{\mathrm{inv}}:=\tfrac{1}{\mu}X^{\top}$—a scaled transpose acting as an approximate inverse—yields a near one‑hot vector on any BB‑sized block); and (c) the leakage patterns of two different rows have small inner product on any such block (small “cross‑leakage”). Given these conditions, we synthesize reference one‑hot weight matrices for keys/queries on each block of size $B$, and realize them in model space by multiplying with $X_{\mathrm{inv}}$. With sparse Bernoulli or Rademacher signatures of width $d_{k}=\Theta(\log m)$ and a fixed dot‑product threshold, each head cleanly separates the true target within its block; max‑pooling over $h\asymp m/B$ heads then recognizes all edges w.h.p., using a total key budget $D_{K}=\Theta\!\big((m/B)\log m\big)$.
Informally, our assumptions are superposition‑style: many items are encoded in a shared low‑dimensional space so that a single linear decoder approximately recovers each one‑hot identity while keeping interference within any block of at most $B$ items uniformly small. Technically, this is a block‑wise low‑coherence (JL/RIP‑style) requirement on $XX^{\top}$, not a general superposition claim. This framework recovers Construction II as a special case: for Gaussian unit–norm embeddings (Cor. B.6) one may take $B=\Theta(d_{\text{model}})$, giving $D_{K}=\Theta\!\big(m\log m/d_{\text{model}}\big)$; for sparse random binary compressive embeddings with $p_{B}=\Theta(\log m/d_{\text{model}})$ (Cor. B.7), one can take $B=\Theta(d_{\text{model}}/\log m)$, giving $D_{K}=\Theta\!\big(m\log^{2}m/d_{\text{model}}\big)$. Full algorithms, constants, and proofs appear in Appendix B.
Construction IV: General Graphs (Summary).
We also extend the compressive permutation scheme (Construction II) to any directed graph $G=(V,E)$ with $|V|=m$ and $|E|=m^{\prime}$. First, we pack $E$ into disjoint matchings of size at most $d_{\text{model}}$ by edge‑coloring the bipartite incidence graph (Kőnig) and batching: this yields $H=\Big\lceil\tfrac{m^{\prime}}{d_{\text{model}}}\Big\rceil+\Delta$ heads, where $\Delta=\max\{\Delta_{\mathrm{out}},\Delta_{\mathrm{in}}\}$. Head $k$ operates on one matching $M_{k}$ (a partial permutation $\pi_{k}:V_{k}\!\to\!T_{k}$) and reuses the Construction II machinery: shared Rademacher signatures, approximate de‑embedding via $X^{\top}$ under Gaussian unit‑norm embeddings, per‑head dimension $d_{k}=C\log m$, and a global threshold $\tau=\tfrac{1}{2}d_{k}$. Exactly one head contains each true edge, while non‑edges lie below $\tau$ in every head; with $d_{\text{model}}\!\geq\!c_{0}\log m$, concentration and a union bound give uniform separation $S^{(k)}_{i,\pi_{k}(i)}>\tau>S^{(k)}_{ij}$ w.h.p., so max‑pooling recovers $E$ and inherits context‑robustness (Lemma 4.1). The total key budget is
$D_{K}\;=\;O\!\left(\frac{m^{\prime}\log m}{d_{\text{model}}}\;+\;\Delta\log m\right).$
Under the mild skew condition $\Delta\leq\frac{m^{\prime}}{d_{\text{model}}}$ (equivalently $\Delta/d_{\mathrm{avg}}\leq m/d_{\text{model}}$ with $d_{\mathrm{avg}}=m^{\prime}/m$), we can take $H=\lceil m^{\prime}/d_{\text{model}}\rceil$ and obtain the tighter $D_{K}=\Theta\!\left(\frac{m^{\prime}\log m^{\prime}}{d_{\text{model}}}\right),$ matching the information‑theoretic lower bound up to constants for all but the densest graphs. Full algorithms, constants, and proofs are deferred to Appendix B.

### v3

@@ 3 Modeling the Self-Attention Mechanism
We model the key–query (QK) computation of a single self-attention layer for RGR, retaining permutation symmetry and parameter sharing while omitting the OV pathway (justified below). The input is an ordered context of distinct vertices of length $\ell\!\leq\!m$, $\mathcal{C}=(v_{i_{1}},\dots,v_{i_{\ell}})$, one vertex per attention unit. Each $v\in V$ is described by a unique embedding $\mathbf{x}_{v}\in\mathbb{R}^{d_{\text{model}}}$. Positional information is not explicitly modeled; if needed, positions can be incorporated by treating $(\text{token},\text{position})$ as distinct vertices.
We here describe the max-over-heads variant of our model; its simplicity lends itself well to more extensive and precise analysis. In Sec. 8 we describe and analyze the more standard version, in which each head applies a softmax over items in $\mathcal{C}$ and the resulting per-head probabilities are summed across heads. In both variants, the benefit of multiple heads arises from noise reduction.
@@
Each attention unit with one head uses the same shared projection matrices $W_{Q},W_{K}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$. For each $v_{i_{p}}\!\in\!\mathcal{C}$, $\mathbf{q}_{i_{p}}=\mathbf{x}_{i_{p}}W_{Q},\mathbf{k}_{i_{p}}=\mathbf{x}_{i_{p}}W_{K}.$ The unnormalized score from source $v_{i_{p}}$ to target $v_{i_{q}}$ is $S_{pq}=\mathbf{q}_{i_{p}}\cdot\mathbf{k}_{i_{q}}^{\top}.$ We declare an edge $(v_{i_{p}},v_{i_{q}})$ present iff $S_{pq}>\tau$ for a global threshold $\tau$. Only pairs inside $\mathcal{C}$ are tested.
Multi-head, max-over-heads version.
With $h$ heads, each head $k$ has $(W_{Q}^{(k)},W_{K}^{(k)})\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ and produces $S_{pq}^{(k)}$. We aggregate by $S_{pq}^{\max}=\max_{k\in\{1,\dots,h\}}S_{pq}^{(k)},$ and decide $(v_{i_{p}},v_{i_{q}})\in E\ \Leftrightarrow\ S_{pq}^{\max}>\tau.$
Remark: why max-over-heads (and why this is not a linear model).
The $\max$ aggregation is a deliberate minimal nonlinearity that preserves the core “competitive routing” behavior of attention while making analysis tractable. It prevents non-target heads from contributing additively to the decision for a given relation, which is a crucial function of softmax as well. If one aggregates scores fully linearly, then multi-heads are equivalent to having a single head (by concatenating the per-head projection matrices Cordonnier et al. (2020a)). [^2] Viewed through an associative-memory lens, purely linear aggregation accumulates cross-term interference, yielding retrieval noise that scales with the number of stored associations relative to key dimension Zhong et al. (2025). Thus, the multi-head advantage requires a nonlinearity to suppress cross-head interference. ’Max-over-heads’ is the cleanest theoretical abstraction of this suppression, and we also replicate key phenomena with standard scaled-softmax (Sec. 8, Sec. 7.1), to show our conclusions are not an artifact of this simpler model.
@@
A construction for RGR maps a graph $G=(V,E)$ to weights $\{(W_{Q}^{(k)},W_{K}^{(k)})\}_{k=1}^{h}$ and a threshold $\tau$ that realize the correct edge decisions for all contexts $\mathcal{C}$, regardless of length $\ell$. We measure complexity by the total key dimension $D_{K}\;=\;h\,d_{k},$ since both compute and parameter footprint for the QK channel scale with the width of the concatenated projections (see Appendix C). [^3] All statements remain (up to a factor of two) with $D_{Q}+D_{K}$. Our goal is to minimize $D_{K}$ over a graph family $\mathcal{G}_{m,m^{\prime}}$ for a given $d_{\text{model}}$.
Analyzing the QK channel in isolation.
Since RGR asks where a source should connect, the key–query computation is the gating step: a correct edge can dominate only if the QK channel already separates the true neighbor from in-context distractors. The OV pathway then acts downstream of this routing decision—reweighting and propagating what QK has selected—so OV can amplify signal but cannot reliably fix systematic mis-routing. [^4] Our model is scoped to a single self-attention layer; multi-layer iterative routing is outside our abstraction. This motivates omitting the OV channel in our model and treating $D_{K}$ as the relevant budget for capacity. This separation between QK (“where”) and OV (“what”) is also supported by recent mechanistic analyses of attention heads Kamath et al. (2025); Franco and Crovella (2025). Finally, our experiments show that the conclusions from this abstraction are robust: when we add back a value channel and when we move to a full Transformer block with realistic frozen embeddings, the same core phenomena predicted by the QK analysis continue to govern performance (Sec. 7.2, Sec. 7.3).
4 Explicit Constructions for RGR
We give QK constructions in the max-over-heads model of Section 3, yielding upper bounds on the key budget $D_{K}=h\,d_{k}$ to solve RGR. This provides a concrete measure of the self-attention mechanism’s capacity for this task. We first sketch the one-hot embedding, permutation graph case (details in App. B), then present our main construction for compressive embeddings, which achieves $D_{K}=\Theta\!\big(\tfrac{m\log m}{d_{\text{model}}}\big)$ for permutation graphs. In App. B, we show how to generalize these results to arbitrary graphs and general embeddings in the max-over-heads model, and in Sec. 8 to permutation graphs in the softmax model. Throughout, $i$ indexes the source and $j$ the target; keys are tied to targets and queries are tied to sources.
Construction I: one-hot permutation graphs.
Let $G$ be a permutation graph on $m$ items with edges $(i,\pi(i))$, and let $\mathbf{x}_{i}=\mathbf{e}_{i}\in\mathbb{R}^{m}$ (so $d_{\text{model}}=m$). Draw signatures $W_{\mathrm{sig}}\in\{\pm 1\}^{m\times d_{k}}$ with i.i.d. Rademacher entries and let $\mathbf{w}_{j}$ be the $j$-th row. Use one head ($h=1$) and set $\mathbf{k}_{j}=\mathbf{w}_{j}$ and $\mathbf{q}_{i}=\mathbf{w}_{\pi(i)}$. Then $S_{i,\pi(i)}=\langle\mathbf{w}_{\pi(i)},\mathbf{w}_{\pi(i)}\rangle=d_{k}$, while for $j\neq\pi(i)$, Bernstein’s inequality shows $S_{ij}=\langle\mathbf{w}_{\pi(i)},\mathbf{w}_{j}\rangle=O(\sqrt{d_{k}})$ w.h.p. Hence choosing $d_{k}=\Theta(\log m)$ and threshold $\tau=d_{k}/2$ yields simultaneous separation over all pairs $(i,j)$ by a union bound, so a single head recovers all edges w.h.p. Details are in App. B. The following shows this implies correctness over all contexts:
Monotonicity under context restriction.
If $S^{\max}_{i,\pi(i)}>\tau$ and $S^{\max}_{ij}<\tau$ for all $j\neq\pi(i)$ over the full vertex set $V$, then the same inequalities hold for any context $\mathcal{C}\subseteq V$ since restricting from $V$ to $\mathcal{C}$ only removes distractor targets.
@@ 4 Explicit Constructions for RGR
We now extend the permutation case to the compressive regime $d_{\text{model}}\ll m$ under a Gaussian unit‑norm embedding: [^5] Random spherical codes let us obtain clean dot‑product thresholds in our separation arguments, but we generalize this to arbitrary embeddings in App. B. This cosine‑geometry is also standard and effective in practice: many systems explicitly constrain features to a hypersphere (e.g., NormFace and ArcFace in face recognition; Spherical Text Embedding in NLP; spherical objectives in metric learning). See also Wang et al. (2017); Deng et al. (2019); Meng et al. (2019); Zhang et al. (2020). Each item $v_{i}$ is embedded as a fixed vector $\mathbf{x}_{i}\in\mathbb{R}^{d_{\text{model}}}$ drawn i.i.d. as $\tilde{\mathbf{x}}_{i}\sim\mathcal{N}(0,I/d_{\text{model}})$ and then $L_{2}$‑normalized, i.e., $\mathbf{x}_{i}=\tilde{\mathbf{x}}_{i}/\|\tilde{\mathbf{x}}_{i}\|_{2}$. Write $X\in\mathbb{R}^{m\times d_{\text{model}}}$ for the matrix with ii‑th row $\mathbf{x}_{i}^{\top}$. Given such an embedding and permutation $\pi$, our goal is to construct attention parameters that recognize $G$.
Multi‑Head Algorithmic Construction.
The fundamental challenge with embeddings is that the input $\mathbf{x}_{i}$ is a superposed representation of the node’s identity. Our construction first approximately inverts the embedding process, projecting the $d_{\text{model}}$-dimensional vector $\mathbf{x}_{i}$ back into the $m$-dimensional one-hot space using the transpose of the embedding matrix. We then apply the logic from the one-hot case. However, doing this with a single head yields too much noise due to the inversion being only approximate. We mitigate this noise by using multiple attention heads, where each is responsible for recognizing the outgoing edges from a disjoint subset of sources. This results in smaller individual heads, and thus less noise. For simplicity, we assume $d_{\text{model}}\mid m$ so that $h=m/d_{\text{model}}$; all bounds and proofs extend to the more general case. The proof of the following theorem appears in Appendix B, where we also show how to extend these results to more general embeddings and graphs.
Algorithm 1 Permutation Graphs with $d_{\mathrm{model}}<m$
@@
2:Parameters: $h=\frac{m}{d_{\text{model}}}$; $d_{k}=C\log m$ for sufficiently large constant $C$.
3:Set Threshold: $\tau=\frac{1}{2}d_{k}$.
4:Partition sources and targets. Split $V$ into $h$ disjoint blocks $V_{1},\dots,V_{h}$ of size $|V_{k}|=d_{\text{model}}$. For each head $k$, define its target set $T_{k}:=\pi(V_{k})=\{\pi(s):s\in V_{k}\}$. $\pi$ is a bijection, so $\{T_{k}\}_{k=1}^{h}$ partition $V$. Head $k$ is responsible for sources in $V_{k}$ and targets in $T_{k}$.
5:Random signatures: Draw $W_{\mathrm{sig}}\in\{\pm 1\}^{m\times d_{k}}$ with i.i.d. Rademacher entries; let $\mathbf{w}_{j}$ be its jj‑th row.
6:Ideal one‑hot‑space templates (for each head $k$):
7:   $\textbf{Query Matrix:}\;W^{\prime}_{Q,(k)}\in\mathbb{R}^{m\times d_{k}}$ with row $i$ equal to $\mathbf{w}_{\pi(i)}$ if $i\in V_{k}$, and $\mathbf{0}$ otherwise.
8:   $\textbf{Key Matrix:}\;W^{\prime}_{K,(k)}\in\mathbb{R}^{m\times d_{k}}$ with row $j$ equal to $\mathbf{w}_{j}$ if $j\in T_{k}$, and $\mathbf{0}$ otherwise.
9:Project back to model space (approximate de‑embedding):
@@
Theorem 4.1 (Multi-head recognition under Gaussian unit-norm embeddings, max-over-heads).
@@
Assume Gaussian unit-norm embeddings with $d_{\mathrm{model}}\geq c_{0}\log m$ for a sufficiently large absolute constant $c_{0}$. Let $h=\frac{m}{d_{\mathrm{model}}}$ heads, per-head dimension $d_{k}=C\log m$ for a sufficiently large absolute constant $C$, and threshold $\tau=\tfrac{1}{2}d_{k}$. Construct $\{(W_{Q}^{(k)},W_{K}^{(k)})\}_{k=1}^{h}$ as in Algorithm 1 and let $k(i)$ denote the unique head index such that $i\in V_{k(i)}$. Then with probability at least $1-m^{-3}$ over the draw of $(X,W_{\mathrm{sig}})$, simultaneously for all $i\in V$:
$S^{(k(i))}_{i,\pi(i)}>\tau\qquad\text{and}\qquad\max_{k\in[h]}\max_{j\neq\pi(i)}S^{(k)}_{ij}<\tau.$
@@
Consequently, $\forall j\neq\pi(i),S^{\max}_{i,\pi(i)}>\tau>S^{\max}_{ij},$ so max-over-heads recovers all edges, and the total key budget satisfies $D_{K}=h\,d_{k}=\Theta\!\Big(\frac{m\log m}{d_{\mathrm{model}}}\Big).$
Consequence.
By Theorem 4.1 we have global separation under max-over-heads, and by monotonicity under context restriction the same parameters recognize $E|_{\mathcal{C}}$ for every context $\mathcal{C}\subseteq V$ and every context length. The total key budget matches our lower bound for permutation graphs up to constant factors. Analogous separation-based arguments translate to softmax; see Sec. 8.